# softMathematical software

### v1

@@ 3.1. Self-adhesion
Proof 3.5.
Extend the bijection $\varphi:N\to M$ by the restriction of $\varphi_{-1}$ to $M\setminus N$ to get a permutation $\psi$ of $NM$. Observe that $\psi({\mathcal{E}})={\mathcal{E}}$. Because $\mathfrak{f}$ is closed under copying, by arguments in Corollary 1, the $\mathfrak{f}$-closure over $NM$ commutes with $\psi$ and one even has $-~\psi(\widetilde{{\mathcal{E}}})=\widetilde{{\mathcal{E}}}$. It follows from the construction and the fact that $\mathfrak{f}$ is closed under marginalization that ${\mathcal{M}}\subseteq{\mathcal{R}}\in\mathfrak{f}(N)$. Using that $-~\widetilde{{\mathcal{E}}}$ is fixed by $\psi$ and the fact that copying commutes with marginalization, one has
@@
Given ${\mathcal{R}}^{\prime}\in\mathfrak{f}(N)$ which is self-adhesive at $L$, and satisfies ${\mathcal{M}}\subseteq{\mathcal{R}}^{\prime}$, it is easy to observe that the respective adhesion ${\mathcal{E}}^{\prime}\in\mathfrak{f}(NM)$ satifies ${\mathcal{E}}\subseteq{\mathcal{E}}^{\prime}$, which yields $\widetilde{{\mathcal{E}}}\subseteq{\mathcal{E}}^{\prime}$ and, hence, ${\mathcal{R}}\subseteq{\mathcal{R}}^{\prime}$. Therefore, one indeed has ${\mathcal{R}}=\mathfrak{f}^{\diamond}({\mathcal{M}}|L)$.
The next example illustrates the method from Lemma 3.4.
@@
Figure 2. The bijection $\varphi:N\to M$ considered in Example 3.6.
This was an example of self-adhesivity at a set. On the other hand, Example 5.1 (in Section 5.3) shows that it may happen that $\mathfrak{sg}^{\diamond}({\mathcal{M}}|L)\neq{\mathcal{M}}$.
@@
3.2. Lifting and tight replication
This section is to show that, if a frame is closed under appropriate operations then the involved models are self-adhesive at sets of particular cardinality.
Definition 3.9 (lifting, tight replication).
  • Let $N\subseteq O$ be two variable sets in inclusion and ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$. The lift of ${\mathcal{M}}$ to $O$ is the following model over $O$:
@@ 3.2. Lifting and tight replication
Lifting complies with marginalization in the sense that, for every ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$ and $N\subseteq O$, one has $({\mathcal{M}}_{N\nearrow O})^{\downarrow N}={\mathcal{M}}$. Analogously, the tight replication also complies with marginalization in the sense that $({\mathcal{M}}_{v\|u})^{\downarrow N}={\mathcal{M}}$.
Figure 3. The mapping $\rho:NM\to N$ considered in Example 3.10.
@@
Put $N:=\{a,b\}$, $O:=\{a,b,c\}$, ${\mathcal{M}}:=\emptyset$, and :=${\mathcal{R}}:=\{\,ab|\emptyset\,\}$. Then ${\mathcal{M}}_{N\nearrow O}=\{\,ac|\emptyset,\,ac|b,\,bc|\emptyset,\,bc|a\,\}$. The interpretation of the lifting is that each element of $O\setminus N$ (in this case, only the variable $c$) becomes independent of the rest. Thus, in case of ${\mathcal{R}}$, one gets the full model: ${\mathcal{R}}_{N\nearrow O}={\mathcal{K}}(O)$.
The interpretation of the tight replication is that one extends the model by introducing a new variable $v$ which is an identical copy of one of the original variables, namely of $u$. To illustrate this concept consider the situation when one has $N:=\{a,b,c\}$, $M:=\{b,c,d\}$, $L:=\{b,c\}$, $u:=a$ and $v:=d$. Define $\rho(a):=a$, $\rho(b):=b$, $\rho(c):=c$, $\rho(d):=a$ (see Figure 3 for illustration) and put ${\mathcal{M}}:=\emptyset$ and :=${\mathcal{R}}:=\{\,ab|c\,\}$. Then
@@
Remark on terminology.   Operations with supermodular set functions, which are analogous to those in Definition 3.9, were discussed in [30, § 7.1 and § 10.2]. Moreover, the concept of tight replication of CI models is a counterpart of an analogous notion of parallel extension of matroids [25, § 7.2] and of polymatroids [23, § 2]. Nevertheless, the parallel extension of polymatroids has two slightly different reflections in the context of induced CI models. In the case that the considered polymatroid $h$ is tight [8, § 1.3] at the variable $u\in N$, which means that $h(N)=h(N\setminus u)$, it leads to our second concept in Definition 3.9. If, however, $h$ is not tight at $u$ then it leads to the concept of parallel extension of CI models introduced in 2004 by Matúš [22, end of § 4]. The extension we need for our purposes is the one that corresponds to a tight polymatroid. Therefore, to avoid a terminological clash, we have chosen a different term than parallel extension. Our terminology also reflects the crucial role of tightness.
Lemma 3.11.
Proof 3.12.
[…]
Given a semi-graphoid ${\mathcal{M}}\in\mathfrak{sg}(N)$ over $N$ and $N\subseteq O$, it is straightforward to observe that ${\mathcal{M}}_{N\nearrow O}$ is a semi-graphoid over $O$.
@@
Proof 3.14.
Consider variable sets $N,M$ assumed in the definition of tight replication and keep the corresponding notational conventions from Definition 3.9. Assume that ${\mathcal{M}}\in\mathfrak{st}(N)$ is induced by a supermodular function $r:{\mathcal{P}}(N)\to{\mathbb{R}}$. Slight modification $\hat{r}$ of $r$ defined by
$\hat{r}(S)~{}:=~{}\left\{\begin{array}[]{cl}r(S)-r(N)+r(L)&\mbox{if $u\in S$,}\\ r(S)&\mbox{if $u\not\in S$,}\\ \end{array}\right.\qquad\mbox{for $S\subseteq N$,}$
is another supermodular function on ${\mathcal{P}}(N)$ which also induces ${\mathcal{M}}$ but, moreover, satisfies $\hat{r}(N)=\hat{r}(L)$. The next step is to define
$m(S)~{}:=~{}\hat{r}(\,(S\cap N)\cup\rho(S\cap M)\,)\qquad\mbox{for~{} $S\subseteq NM$}\,.$
For any $ij|K\in{\mathcal{K}}(NM)$ one has $\Delta m(ij|K)=\Delta\hat{r}\,\langle\rho(i),\rho(j)|\rho(K)\rangle$. The latter expression is non-negative in case $\rho(ij)|\rho(K)\in{\mathcal{K}}(N)$ by supermodularity of $\hat{r}$. Moreover, the expression vanishes if $\rho(ij)\cap\rho(K)\neq\emptyset$. In the remaining case $\rho(i)=u=\rho(j)$ and $C:=\rho(K)\subseteq L$ one has
$\Delta\hat{r}\,\langle\rho(i),\rho(j)|\rho(K)\rangle=\Delta\hat{r}\,\langle u,u|C\rangle=\hat{r}(C)-\hat{r}(uC)=\Delta\hat{r}\,\langle u,L\setminus C|C\rangle\geq 0\,,$
by supermodularity of $\hat{r}$. Hence, $m$ is supermodular and it follows from the definition of tight replication that ${\mathcal{M}}^{m}={\mathcal{M}}_{v\|u}$.
@@
One can distinguish several cases determined by the mutual position of variables $u$ and $v$ relative to $i,j,\ell,K$. If $uv\setminus ij\ell K\neq\emptyset$ then the conclusion follows from the facts that ${\mathcal{E}}^{\downarrow N}={\mathcal{M}}$ and ${\mathcal{E}}^{\downarrow M}$ is a copy of ${\mathcal{M}}$. The same principle applies if $u,v\in K$. In cases $|uv\cap K|=1=|uv\cap ij\ell|$ the first convention in the definition of ${\mathcal{M}}_{v\|u}$ is applied. This also extends to the case $uv=j\ell$. In the cases $uv=ij$ and $uv=i\ell$ one additionally applies the second convention in the definition of ${\mathcal{M}}_{v\|u}$ together with the assumption that ${\mathcal{M}}$ is a semi-graphoid.
The lifting operation is related to adhesions at small sets.
@@
Let $\mathfrak{f}\subseteq\mathfrak{sg}$ be a CI frame which is closed under marginalization, intersection, and lifting. Then, for every ${\mathcal{M}}\in\mathfrak{f}(N)$ and ${\mathcal{R}}\in\mathfrak{f}(M)$ such that $|N\cap\/M|\leq 1$, an adhesion ${\mathcal{E}}\in\mathfrak{f}(NM)$ of ${\mathcal{M}}$ and ${\mathcal{R}}$ exists.
Proof 3.16.
Put ^$\hat{{\mathcal{M}}}:={\mathcal{M}}_{N\nearrow NM}$ and observe that $(N\setminus M)\perp\!\!\!\perp(M\setminus N)\,|\,(N\cap\/M)\,\,[\hat{{\mathcal{M}}}]$ and ^$\hat{{\mathcal{M}}}^{\downarrow N}={\mathcal{M}}$. Moreover, $|N\cap\/M|\leq 1$ implies ^$\hat{{\mathcal{M}}}^{\downarrow M}={\mathcal{K}}(M)$ and the fact ^$\hat{{\mathcal{M}}}\in\mathfrak{f}(NM)$ follows from the assumption that $\mathfrak{f}$ is closed under lifting. Putting ^$\hat{{\mathcal{R}}}:={\mathcal{R}}_{M\nearrow NM}\in\mathfrak{f}(NM)$ gives $(N\setminus M)\perp\!\!\!\perp(M\setminus N)\,|\,(N\cap\/M)\,\,[\hat{{\mathcal{R}}}]$ with ^$\hat{{\mathcal{R}}}^{\downarrow N}={\mathcal{K}}(N)$ and ^$\hat{{\mathcal{R}}}^{\downarrow M}={\mathcal{R}}$. Since $\mathfrak{f}$ is closed under intersection, $-^{\mathcal{E}}:=\hat{{\mathcal{M}}}\cap\hat{{\mathcal{R}}}$ belongs to $\mathfrak{f}(NM)$ and is an adhesion of ${\mathcal{M}}$ and ${\mathcal{R}}$.
Corollary 4.
Let $\mathfrak{f}\subseteq\mathfrak{sg}$ be a CI frame which is closed under marginalization, copying, intersection, and lifting. Then, for any $N$, any model ${\mathcal{M}}\in\mathfrak{f}(N)$ is self-adhesive at any set of cardinality 00 or $1$.
Proof 3.17.
Apply Lemma 3.15 to ${\mathcal{M}}$ and ${\mathcal{R}}:=\varphi({\mathcal{M}})$ for a bijection $\varphi:N\to M$.
The tight replication is related to self-adhesions at large sets.
Lemma 3.18.
Let $\mathfrak{f}\subseteq\mathfrak{sg}$ be a CI frame which is closed under marginalization, copying, and tight replication. Then, for any variable set $N$ of cardinality $n\geq 1$, any model ${\mathcal{M}}\in\mathfrak{f}(N)$ is self-adhesive at any set of cardinality $n-1$.
Proof 3.19.
Consider ${\mathcal{M}}\in\mathfrak{f}(N)$, a subset $L\subseteq N$ with $|L|=n-1$, and a bijection $\varphi:N\to M$ such that $N\cap\/M=L$ and $\varphi_{\,|L}=\mbox{\rm id}_{L}$. Put $u:=N\setminus L$ and $v:=M\setminus L$. By the assumption, ${\mathcal{E}}:={\mathcal{M}}_{v\|u}$ belongs to $\mathfrak{f}(NM)$. The definition of tight replication allows one to observe that ${\mathcal{E}}^{\downarrow N}={\mathcal{M}}$, ${\mathcal{E}}^{\downarrow M}=\varphi({\mathcal{M}})$, and $u\perp\!\!\!\perp v\,|\,L\,\,[{\mathcal{E}}]$. Hence, ${\mathcal{E}}$ is an adhesion of ${\mathcal{M}}$ and $\varphi({\mathcal{M}})$.
Corollary 5.
If a CI frame $\mathfrak{f}\subseteq\mathfrak{sg}$ is closed under marginalization, copying, intersection, lifting, and tight replication and $|N|\leq 3$ then $\mathfrak{f}^{\diamond}(N)=\mathfrak{f}(N)$. In particular, $\mathfrak{sg}^{\diamond}(N)=\mathfrak{sg}(N)$ and $\mathfrak{st}^{\diamond}(N)=\mathfrak{st}(N)$ in case $|N|\leq 3$.
Proof 3.20.
Combine Corollary 4, Lemma 3.18 and the fact that any ${\mathcal{M}}\in\mathfrak{f}(N)$ is self-adhesive at $L:=N$. Use Lemmas 2.6, 2.11, 3.11, and 3.13 for the second part.
The second part of Corollary 5 is not surprising in light of Lemma 3.7 since $\mathfrak{sg}(N)=\mathfrak{st}(N)=\mathfrak{pr}(N)$ if $|N|\leq 3$. On the other hand, the observations in Section 3.2 allow one to simplify testing self-adhesive models in case $|N|\geq 4$.
4. Iterations of the self-adhesion operator

### v3

@@ 4.1. Self-adhesion
Proof.
Extend the bijection $\varphi:N\to M$ by the restriction of its own inversion $\varphi_{-1}$ to $M\setminus N$ to get a permutation $\psi$ of $NM$. Observe that $\psi({\mathcal{E}})={\mathcal{E}}$. Because $\mathfrak{f}$ is closed under copying, by arguments in Observation 2.7, the $\mathfrak{f}$-closure over $NM$ commutes with $\psi$ and one even has $\psi(\widetilde{{\mathcal{E}}})=\widetilde{{\mathcal{E}}}$. It follows from the construction and the fact that $\mathfrak{f}$ is closed under marginalization that ${\mathcal{M}}\subseteq{\mathcal{R}}\in\mathfrak{f}(N)$. Using that $\widetilde{{\mathcal{E}}}$ is fixed by $\psi$ and the fact that copying commutes with marginalization, one has
@@
Given ${\mathcal{R}}^{\prime}\in\mathfrak{f}(N)$ which is self-adhesive at $L$, and satisfies ${\mathcal{M}}\subseteq{\mathcal{R}}^{\prime}$, it is easy to observe that the respective adhesion ${\mathcal{E}}^{\prime}\in\mathfrak{f}(NM)$ satisfies ${\mathcal{E}}\subseteq{\mathcal{E}}^{\prime}$, which yields $\widetilde{{\mathcal{E}}}\subseteq{\mathcal{E}}^{\prime}$ and, hence, ${\mathcal{R}}\subseteq{\mathcal{R}}^{\prime}$. Therefore, one indeed has ${\mathcal{R}}=\mathfrak{f}^{\diamond}({\mathcal{M}}|L)$. ∎
Note that the $\mathfrak{f}$-closure is applied only once and it is over a variable set of cardinality $|N|+|L|$. The next example illustrates the method from Lemma 4.5.
@@
This was a positive example of self-adhesivity at a set. On the other hand, Example 7.1 (in Section 7.3), obtained as a result of our computations, shows that it may happen that $\mathfrak{sg}^{\diamond}({\mathcal{M}}|L)\neq{\mathcal{M}}$.
@@
4.2. Self-adhesivity at small sets: lifting
In next two subsections we show that, if a CI frame is closed under appropriate extending algebraic operations, then the involved models are automatically self-adhesive at sets of particular cardinality.
The first such operation, called lifting, ensures the self-adhesivity at the empty set and singletons. It formalizes the act of adding completely independent variables. Specifically, in case of a CI model ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$ and a single (outside) variable $o\not\in N$, the respective (one-variable) lift is the following model over $No$ :
${\mathcal{M}}\,\cup\,\{\,ij|Ko\,:\ ij|K\in{\mathcal{M}}\,\}\,\cup\,\{\,io|K\,:\ i\in N,\,K\subseteq N\setminus i\,\}\,.$
Such one-element (independent) additions commute each other and allow one to introduce the addition of a set of independent variables. Nonetheless, for technical reasons, we prefer the following condensed definition, which is equivalent to the result of repeated one-variable additions.
Definition 4.8 (lifting).
Let $N\subseteq O$ be variable sets in inclusion and ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$. The lift of ${\mathcal{M}}$ to $O$ is the following model over $O$:
@@ 4.2. Self-adhesivity at small sets: lifting
Note that lifting complies with marginalization in the sense that, for every ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$ and $N\subseteq O$, one has $({\mathcal{M}}_{N\nearrow O})^{\downarrow N}={\mathcal{M}}$.
Example 4.9.
Put $N:=\{a,b\}$, $O:=\{a,b,c\}$, ${\mathcal{M}}:=\emptyset$, and :=${\mathcal{R}}:=\{\,ab|\emptyset\,\}$ and interpret both ${\mathcal{M}}$ and ${\mathcal{R}}$ as CI models over $N$. Then ${\mathcal{M}}_{N\nearrow O}=\{\,ac|\emptyset,\,ac|b,\,bc|\emptyset,\,bc|a\,\}$. Note that the variable $c$ became independent of the rest. In the other case of ${\mathcal{R}}$, one gets the full model: ${\mathcal{R}}_{N\nearrow O}={\mathcal{K}}(O)$.
Lemma 4.10.
Proof.
[…]
Given a semi-graphoid ${\mathcal{M}}\in\mathfrak{sg}(N)$ over $N$ and $N\subseteq O$, it is straightforward to observe that ${\mathcal{M}}_{N\nearrow O}$ is a semi-graphoid over $O$. ∎
The lifting operation is related to adhesions of CI models at small sets.
Lemma 4.11.
Let $\mathfrak{f}$ be a CI frame which is closed under marginalization, intersection, and lifting. Then, for every ${\mathcal{M}}\in\mathfrak{f}(N)$ and ${\mathcal{R}}\in\mathfrak{f}(M)$ such that $|N\cap\/M|\leq 1$, an adhesion ${\mathcal{E}}\in\mathfrak{f}(NM)$ of ${\mathcal{M}}$ and ${\mathcal{R}}$ exists.
Proof.
Put ^$\widehat{{\mathcal{M}}}:={\mathcal{M}}_{N\nearrow NM}$ and observe that $(N\setminus M)\perp\!\!\!\perp(M\setminus N)\,|\,(N\cap\/M)\,\,[\widehat{{\mathcal{M}}}]$ and ^$\widehat{{\mathcal{M}}}^{\downarrow N}={\mathcal{M}}$. Moreover, $|N\cap\/M|\leq 1$ implies ^$\widehat{{\mathcal{M}}}^{\downarrow M}={\mathcal{K}}(M)$ and the fact ^$\widehat{{\mathcal{M}}}\in\mathfrak{f}(NM)$ follows from the assumption that $\mathfrak{f}$ is closed under lifting. Putting ^$\widehat{{\mathcal{R}}}:={\mathcal{R}}_{M\nearrow NM}\in\mathfrak{f}(NM)$ gives $(N\setminus M)\perp\!\!\!\perp(M\setminus N)\,|\,(N\cap\/M)\,\,[\widehat{{\mathcal{R}}}]$ with ^$\widehat{{\mathcal{R}}}^{\downarrow N}={\mathcal{K}}(N)$ and ^$\widehat{{\mathcal{R}}}^{\downarrow M}={\mathcal{R}}$. Since $\mathfrak{f}$ is closed under intersection, $^{\mathcal{E}}:=\widehat{{\mathcal{M}}}\cap\widehat{{\mathcal{R}}}$ belongs to $\mathfrak{f}(NM)$ and, moreover, by the above facts, ${\mathcal{E}}$ is an adhesion of ${\mathcal{M}}$ and ${\mathcal{R}}$. ∎
Corollary 4.12.
Let $\mathfrak{f}$ be a CI frame which is closed under copying, marginalization, intersection, and lifting. Then, for any $N$, any model ${\mathcal{M}}\in\mathfrak{f}(N)$ is self-adhesive at every set of cardinality 00 or $1$.
Proof.
Apply Lemma 4.11 to ${\mathcal{M}}$ and ${\mathcal{R}}:=\varphi({\mathcal{M}})$ for the appropriate bijection $\varphi:N\to M$. ∎
4.3. Self-adhesivity at co-singletons: tight replication
In this subsection we discuss a particular extending algebraic operation which ensures that, if a CI frame $\mathfrak{f}$ is closed under this operation, then every ${\mathcal{M}}\in\mathfrak{f}(N)$ is self-adhesive at co-singletons, that is, at the complements of singletons within the variable set $N$. This operation is called tight replication and formalizes the act of adding one identical copy of a fixed variable in $N$. This (technically complex) operation was inspired by an analogous (poly-)matroidal operation of parallel extension.
Let us first remind what is our goal and re-formulate the condition of self-adhesivity at a co-singleton (we wish to ensure). Consider ${\mathcal{M}}\in\mathfrak{f}(N)$, fix a variable $u\in N$ and an outside variable $v\not\in N$. Consider the co-singleton $L:=N\setminus u$, denote $M:=Lv$, and introduce $\psi:N\to M$ which is identical on $L$ and $\psi(u)=v$. By Definition 4.3, our goal is to verify the existence of
${\mathcal{E}}\in\mathfrak{f}(NM)$   which satisfies   ${\mathcal{E}}^{\downarrow N}={\mathcal{M}}$, ${\mathcal{E}}^{\downarrow M}=\psi({\mathcal{M}})$, and $u\perp\!\!\!\perp v\,|\,L\,\,[{\mathcal{E}}]$.
One of possible ways is to define ${\mathcal{E}}$ as a particular extension of ${\mathcal{M}}$ in which $v$ is the “identical" copy of $u$. To this end an auxiliary mapping $\rho:NM\to N$ is used, which an extension of the inverse $\psi_{-1}:M\to N$.
Definition 4.13 (tight replication).
@@ 4.3. Self-adhesivity at co-singletons: tight replication
Note that tight replication also complies with marginalization in the sense that ${({\mathcal{M}}_{v\|u})^{\downarrow N}={\mathcal{M}}}$.
Figure 3. The auxiliary mapping $\rho:NM\to N$ considered in Example 4.14.
@@
Consider the situation from Definition 4.13 when one has $N:=\{a,b,c\}$, $M:=\{b,c,d\}$, $L:=\{b,c\}$, $u:=a$ and $v:=d$. Define $\rho(a):=a$, $\rho(b):=b$, $\rho(c):=c$, $\rho(d):=a$ (see Figure 3 for illustration), put ${\mathcal{M}}:=\emptyset$ and :=${\mathcal{R}}:=\{\,ab|c\,\}$. Then
@@
Remark on terminology.   Operations with supermodular set functions, which are analogous to those in Definitions 4.8 and 4.13, were discussed in [38, § 7.1 and § 10.2]. Moreover, the concept of tight replication of CI models is a counterpart of an analogous notion of parallel extension of matroids [31, § 7.2] and of polymatroids [29, § 2]. Nevertheless, the parallel extension of polymatroids has two slightly different reflections in the context of induced CI models. In the case that the considered polymatroid $h$ is tight [9, § 1.3] at the variable $u\in N$, which means that $h(N)=h(N\setminus u)$, it leads to our concept from Definition 4.13. If, however, $h$ is not tight at $u$ then it leads to the concept of parallel extension of CI models introduced in 2004 by Matúš [28, end of § 4]. The extension we need for our purposes is the one that corresponds to a tight polymatroid. Therefore, to avoid a terminological clash, we have chosen a different term than parallel extension. Our terminology also reflects the crucial role of tightness.
@@
Proof.
Consider variable sets $N,M$ assumed in the definition of tight replication and keep the corresponding notational conventions from Definition 4.13. Assume that ${\mathcal{M}}\in\mathfrak{st}(N)$ is induced by a supermodular function $r:{\mathcal{P}}(N)\to{\mathbb{R}}$. Slight modification $\widehat{r}$ of $r$ defined by
$\widehat{r}(S)~{}:=~{}\left\{\begin{array}[]{cl}r(S)-r(N)+r(L)&\mbox{if $u\in S$,}\\ r(S)&\mbox{if $u\not\in S$,}\\ \end{array}\right.\qquad\mbox{for $S\subseteq N$,}$
is another supermodular function on ${\mathcal{P}}(N)$ which also induces ${\mathcal{M}}$ but, moreover, satisfies $\widehat{r}(N)=\widehat{r}(L)$. The next step is to define
$m(S)~{}:=~{}\widehat{r}(\,(S\cap N)\cup\rho(S\cap M)\,)\qquad\mbox{for~{} $S\subseteq NM$}\,.$
For any $ij|K\in{\mathcal{K}}(NM)$ one has $\Delta m(ij|K)=\Delta\widehat{r}\,\langle\rho(i),\rho(j)|\rho(K)\rangle$. The latter expression is non-negative in case $\rho(ij)|\rho(K)\in{\mathcal{K}}(N)$ by supermodularity of $\widehat{r}$. Moreover, the expression vanishes if $\rho(ij)\cap\rho(K)\neq\emptyset$. In the remaining case $\rho(i)=u=\rho(j)$ and $C:=\rho(K)\subseteq L$ one has
$\Delta\widehat{r}\,\langle\rho(i),\rho(j)|\rho(K)\rangle=\Delta\widehat{r}\,\langle u,u|C\rangle=\widehat{r}(C)-\widehat{r}(uC)=\Delta\widehat{r}\,\langle u,L\setminus C|C\rangle\geq 0\,,$
by supermodularity of $\widehat{r}$. Hence, $m$ is supermodular and it follows from the definition of tight replication that ${\mathcal{M}}^{m}={\mathcal{M}}_{v\|u}$.
@@
One can distinguish several cases determined by the mutual position of variables $u$ and $v$ relative to $i,j,\ell,K$. If $uv\setminus ij\ell K\neq\emptyset$ then the conclusion follows from the facts that ${\mathcal{E}}^{\downarrow N}={\mathcal{M}}$ and ${\mathcal{E}}^{\downarrow M}$ is a copy of ${\mathcal{M}}$. The same principle applies if $u,v\in K$. In cases $|uv\cap K|=1=|uv\cap ij\ell|$ the first convention in the definition of ${\mathcal{M}}_{v\|u}$ is applied. This also extends to the case $uv=j\ell$. In the cases $uv=ij$ and $uv=i\ell$ one additionally applies the second convention in the definition of ${\mathcal{M}}_{v\|u}$ together with the assumption that ${\mathcal{M}}$ is a semi-graphoid. ∎
Remark. Note in connection with Lemma 4.15 that have not tried to answer the question whether $\mathfrak{pr}$ is closed under tight replication. If it is the case then this certainly requires a technically complicated proof, but the point is that such a result is not needed for our purpose. Indeed, we know, by Lemma 4.7, that $\mathfrak{pr}$ is self-adhesive, and, thus, self-adhesive at co-singletons.
@@
Let $\mathfrak{f}$ be a CI frame which is closed under copying, marginalization, and tight replication. Then, for any variable set $N$ of cardinality $n\geq 1$, any model ${\mathcal{M}}\in\mathfrak{f}(N)$ is self-adhesive at every co-singleton.
Proof.
Consider ${\mathcal{M}}\in\mathfrak{f}(N)$, a subset $L\subseteq N$ with $|L|=n-1$ (= a co-singleton), and a bijection $\psi:N\to M$ such that $N\cap\/M=L$ and $\psi_{\,|L}=\mbox{\rm id}_{L}$. Put $u:=N\setminus L$ and $v:=M\setminus L$. By the assumption, ${\mathcal{E}}:={\mathcal{M}}_{v\|u}$ belongs to $\mathfrak{f}(NM)$. The definition of tight replication allows one to observe that ${\mathcal{E}}^{\downarrow N}={\mathcal{M}}$, ${\mathcal{E}}^{\downarrow M}=\psi({\mathcal{M}})$, and $u\perp\!\!\!\perp v\,|\,L\,\,[{\mathcal{E}}]$. Hence, ${\mathcal{E}}$ is an adhesion of ${\mathcal{M}}$ and $\psi({\mathcal{M}})$. ∎
We now apply the observations from Sections 4.2 and 4.3.
Corollary 4.17.
If a CI frame $\mathfrak{f}$ is closed under copying, marginalization, intersection, lifting, and tight replication and $|N|\leq 3$ then $\mathfrak{f}^{\diamond}(N)=\mathfrak{f}(N)$. In particular, $\mathfrak{sg}^{\diamond}(N)=\mathfrak{sg}(N)$ and $\mathfrak{st}^{\diamond}(N)=\mathfrak{st}(N)$ if $|N|\leq 3$.
Proof.
Combine Corollary 4.12, Lemma 4.16 and the fact that any ${\mathcal{M}}\in\mathfrak{f}(N)$ is self-adhesive at $L:=N$. Use Lemmas 3.1, 3.4, 4.10, and 4.15 for the second part. ∎
The second part of Corollary 4.17 is not surprising in light of Lemma 4.7 since $\mathfrak{pr}(N)=\mathfrak{st}(N)=\mathfrak{sg}(N)$ if $|N|\leq 3$. On the other hand, Corollary 4.12 and Lemma 4.16 allow one to simplify testing self-adhesive models in case of large variable sets $N$ with $|N|\geq 4$.
5. Self-adhesion operator and its iterating
In this section we discuss a self-adhesion operator which assigns a subframe $\mathfrak{f}^{\diamond}$ to an abstract CI frame $\mathfrak{f}$, closed under copying and marginalization. We prove the consistency results for such a construction, present an algorithm to compute the respective $\mathfrak{f}^{\diamond}$-closure, and discuss the iterations of the self-adhesion operator.